OpenStudy (rissyroo118):

a company purchased a new truck for $84,000 on july 1. the truck is estimated to have a useful life of 6 years and a salvage value of $12,000. using the straight line method , determine the amount of depreciation expenses recorded during the first year ending december 31 $12,000 $3,000 $7,000 $6,000

OpenStudy (rissyroo118):

@mathmate

OpenStudy (mrnood):

if they buy it for 84000 and sell at the end for 12000 how much has it cost them IN TOTAL?

OpenStudy (mrnood):

@rissyroo118 Please tell me if you are going to answer - I will go elsewhere if not....

OpenStudy (mathmate):

@rissyroo118 @MrNood is you best bet for this one!

OpenStudy (mrnood):

You will get more help if you show some inclination to help yourself. Please do not expect full answer given to you

OpenStudy (rissyroo118):

im sorry i was busy, im at work doing these problems on here as well as working. i goot busy

OpenStudy (rissyroo118):

@mrnood

OpenStudy (mrnood):

so if they buy it for 84000 and sell at the end for 12000 how much has it cost them IN TOTAL?

OpenStudy (rissyroo118):

84,000-12,000 = 72,000?

OpenStudy (mrnood):

yes and that was spread over 6 years so how much did it cost each year?

OpenStudy (mrnood):

sorry - can't keep waiting hint the period they ask you for is 6 months - 0.5 year so get the cost per year then work out how much that is in 6 months

OpenStudy (rissyroo118):

72,000/6= 12,000

OpenStudy (rissyroo118):

@mrnood

OpenStudy (mrnood):

so get the cost per year then work out how much that is in 6 months

OpenStudy (rissyroo118):

so it would be 12,000 a year , 6/12= 2,000 is what i get i dont know what im doing wrong

OpenStudy (rissyroo118):

wait i did it wrong i think, so i did 12,000/2=6,000 @MrNood so D would be the answer

OpenStudy (mrnood):

Yes - seems like you knew how to do it all along?

OpenStudy (rissyroo118):

took some time to figure out what you wanted me to do, but then i got it and started to understand. i really appreciate your help :)